FEL Upgrade Project Weekly Brief
August 27-31, 2001



 
Highlights:
The FEL Weekly Report returns from a week's hiatus due to last week's International FEL Conference in Darmstadt.  High
average power FEL's, energy recovery schemes, and FEL specific applications were well represented at the conference.

Our readers should note that an article on the recent user applications of the IR Demo FEL was just published in the August
issue of Laser Focus World (including a nice cover photo about the nanotube production by the CWM-Penn State, NASA
team).  The article is available at the following website:  www.optoelectronics-world.com.

Plans are being put together this week for the last run of the IR Demo starting next month and ending with the presently
scheduled shut down on Dec. 1 for the start of upgrade installation activities.

Bids were received this week for fabrication of the Upgrade GW dipole magnets.  They are currently under technical and cost
evaluation.

Management:
At last week's FEL Conference at Darmstadt, the FEL Program at Jefferson Lab was well represented.  Both George Neil and
Steve Benson gave invited talks as follow-up to their winning last year's "FEL Prize".  George discussed the evolution of light
source development over the last 10 years with the recent successes with amplifier FELs (SASE's), high average power
demonstrations with energy recovery schemes, and extensions of the energy recovery concept to next generation x-ray sources.
Steve gave a summary presentation on what we have learned on FEL related physics and technology development from the IR
Demo.  Lia Merminga gave an invited talk on "RF Stability in Energy Recovering FELs" and Ed Gillman gave a presentation in
the FEL User's Workshop on his work on scanning near field microscopy.  Additional presentations by JLab personnel included:
"Mode Distortion Measurements" by S. Benson et al; "Second Harmonic FEL Oscillation" by G. Neil et al; "Design Principles for
a Compact High Average power FEL" by L Merminga and S. Benson.  Several papers were presented by our colleagues from
other institutions on work related to FELs at JLab.  Bill Colson's group at NPS presented two posters on their preliminary
simulations of a 100 kW design using short Raleigh lengths and a step tapered undulator.  Alan Todd at AES gave a poster on the
AES model for high average power FELs based on SRF energy recovered linacs. Conference papers are currently in the review
cycle prior to publication.

With respect to the last run of the IR Demo prior to planned December 1 shutdown, we are collecting data from our internal and
external users to use this time most effectively.  External user requests are due on Sept. 15th.  Prior to the external user run
which begins in October, we will be running the IR Demo for the second phase of the JTO laser damage studies and for machine
development needs for the upgrade.

We were in frequent communication with our Program Officer at AFRL to help complete the last paperwork for the start-up of
the UV FEL project.

We are analyzing end of year budget projections prior to closing out the Phase1 of the FEL Upgrade Project on Sept. 30.

WBS 4 (Injector):
The external corona rings were received and QCed.  The drawings for the cesium channel buses and charge collection ring have
been signed off and submitted to the shop.  Machining is being done on the gun chamber parts and the ball cathode.

Gun's HVPS - The HVPS Tank was opened to provide the Mechanical Engineer and Designer an opportunity to take field
measurements so their new design will mate with the existing Transmission Line and Gun.  The multiplier stacks were also
temporarily assembled for additional measurements to design the new HVPS Tank.

WBS 5(SRF):
We have received a very attractive offer from Toshiba for electropolishing the SRF cavities for the upgrade cryomodule.  A
contract will be pursued in the Phase 2 effort

WBS 6 (RF):
Zone 3 - Corrected door interlock wiring errors this week and made list of changes.  Missing the 5 watt preamps and their power
supply, but they are in-house and scheduled to be installed next week.  The HVPS will be no-load tested early next week.  After
the preamps are installed, the zone will be operated into shorts for the first tests.

Zone 4 - The 8 RF Control Modules and 2 stepper cards are still missing, but AES is working to replace them.

Quarter HVPS - Parts for the HVPS are being ordered and some have already been received.  A red-marked schematic was
turned into Drafting as a starting point document.  The 3 new 100 kW Circulators were tested this week.  They do not meet
specifications completely and we able to test them to only 50 kW, but they appear to be workable.  The vendor is working to help
us adjust the Circulator S11 and S21 bandwidths to be flatter and wider.

WBS 8 (Instrumentation):
Testing is nearly complete on the VME timing/delay board and EPICS software is being developed to support it.  Prototype
boards for the new beamviewer camera/lamp board power supplies have arrived and all components for the beamviewer control
box are on order.  Additional carrier boards have been ordered to be used as new systems are developed.
 
An analog differential driver board is also under development for better monitoring of beam current. The prototype has been
installed and works very well. This reduces the noise on the bunch charge monitor as well as providing ~1 microsecond rise time
to allow fault-free alignment mode operation with macropulses less than 10 microseconds.
 
The gun high voltage tank SF6 pressure transducer & control room readout has been commissioned and is operational.  This had
been on the "list" longer than most.

The first of the new Ion Pump Power Supplies has been shipped back to the lab for testing.

The BAIR web and the task system are continually being upgraded to make them more useful.  A digital Input/Output box is
being built for use in the labs to allow more functionality and control of test equipment.

Drawings completed this week:  F0015 - 32 Channel Sample and Hold Board drawings have been revised to correct the
placement of the connectors.  F0060 - VME Timing Main Board was revised to make a minor change in hole size for the
connector and one of the front test connections.  F0094-98 - Temperature Monitoring Chassis:  Wiring Diagram, Fabrication, and
Assembly Drawings

WBS 9 (Transport):
Dipoles
Optical Chicane Dipoles (GW)
o  Proposals were received and evaluated.  Several items needed clarification from the offerors.
Injector Dipoles (DU/DV)
o  The final round of rework on the Small Magnet's (GV) Assembly Drawing is nearly complete.  They started the final correction of
    the Large Magnetís (GU) details.
Arc Dipoles (GY, GX, GQ)
o  David Douglas analyzed the magnetic model data for the 180 Dipole (GY) from AES.  The result was permission to complete the
    drawings per the magnetic model.  His specifications were met.
o  AES has almost completed their manipulation of the magnetic model of the Reverse Bend (GQ).  They had to add a .001 inch
    step to the edge of the pole, adjust the stepís width and add a taper to get uniform field and integral.  They have yet to adjust the
    gap height and faces to get the correct bulk field and effective length.
o  Design continued on backchecking the detail drawings of all three magnets with the comments and corrections.  They also started
    adjusting the DY drafting model (and hence, automatically, most of the details) to the dimensions of the successful magnetic
    model.
Quadrupoles
3 inch quad (QX)
o  The specification for assembly and the assembly drawing were signed and placed in the procurement cycle.  Work continues on
    cores and coils.
Trim Quad (QT)
o  The 2D detail drawings and statement of work are in final checking
Sextupole (SF)
o  Work slowed as the octupole statement of work and quadrupole assembly and a well-earned vacation took Robin Wineís time.
Octupole (OT)
o  DULY Research had results from their magnetic modeling.  They found that a mushroom cap pole tip of the right shape, placed
    on the recycled corrector coils gets a 2 to 3 % magnet with simple flat truncated ends.  They are continuing their pursuit of the
    1% goal by adding field clamps, nubbins at the pole tip edge and rounding the pole ends. Analysis shows that the next level power
    supplies used for the Trim Quads (20 A, 50 V) will be necessary for these magnets also.  Kevin Jordan will be happy about this.
    Not!
Corrector Dipoles (DB, DJ)
o  Milhous Co., in developing their production process, made a corrector coil that looked good on their second try but it had a short.
o  We received feedback on our concept design of the new mounting bracket for the DB/DJ correctors from the plastic molding
    company.  Those suggestions are being incorporated into a final design.
Beam Line and Vacuum
o  Work continued on the definition of the Optical Chicane chambers and continued to the mounting stand for the GW magnets.
o  The bid package drawing set for the girders in the return leg continues in detailing.
o  We continued engineering work on the LCW distribution.
o  Butch Dillon-Townes is working on finalizing the style of the recirculation dump.  It still is subject to controversy.  George Neil =
    garbage can style.  Dave Douglas = a narrow design.
o  Work started on the injector lines stands and vacuum chamber.  Here also, a dispute about the configuration of the chambers has
    arisen and is being resolved by Butch Dillon-Townes.

WBS 10 (Wiggler):
The Sentron 3-axis probe has been calibrated using an arc dipole magnet.  The probe shows no evidence of a transverse Hall
effect at 3 kG so we should be able to measure all three axes simultaneously.  The zero drift in the probe, though within
specifications, is higher than in data provided by Sentron so we are exploring ways of reducing the drift.  The measurement
assembly is calibrated and ready for installation.  We should be able to start measurements next week.  The corrector coils are
installed on the dispersion magnet.  The only work left on this magnet is to finish wiring up the corrector coils to the junction
blocks.  This magnet will be measured after the wiggler magnets.  The dispersion section magnet vacuum chamber is done.  The
wiggler vacuum chambers are ready for welding on the side tubes.  The chambers should be complete in two weeks.

WBS 11 (Optics):
We've been refining our parts list for the optical cavity chamber which took somewhat longer than anticipated, due to the need to
get some of the parts machined to high tolerance.  We met with Survey and Alignment to resolve some possible interference
issues, and believe we have a solution.  The PO for the linear actuators was awarded.  More fabricated parts for the
ultraviewers arrived, we are awaiting only two more.  We purchased more IR cameras to assist in beam profiling and position
monitoring.

Only one vendor submitted a bid to our drive laser solicitation.  The price is far higher than our estimate, so we are working with
them to understand their price discrepancy.  We are also revisiting alternate scenarios for building our own laser.

Support for operations/experimenters:
There was no operations the week of 8/20 and the first part of 8/27, however, we prepared for further characterization of THz
radiation (by adding a breadboard and better motion control) in the vault.  We also prepared for a series of measurements of the
FEL phase noise (timing jitter) and how it compares vis-a-vis the drive laser.  This included running a low-loss silica optical fiber
between the drive laser enclosure (DLE) and the optics control room (OCR), running RF cabling out of the DLE for remote
measurements.  It also included work by the Accel. Div. RFES group, who fielded measurement equipment, and
hardware/support.

Experiments began yesterday (8/30) and preliminary results will be reported next week.  In order to better understand our IR
Demo mirror distortion measurements, we plan to do some characterization of the measuring device, a Coherent ModeMaster.
To do this, we set up our small carbon dioxide laser in User Lab 2.  We will do the characterization next week,  after it is
certified for safe use.

To support this same experiment, we characterized the plano sides of our calcium fluoride cavity mirrors.  We found the mirrors
had about a half-wave convex (ROC ~ 1 km) figure.  We need to do further measurements to see if this is due to the mounting,
or residual error after the coating.  We also supported the Accel. Div. by measuring the wavefront distortion in some of the
viewports they wish to use for new synchrotron light monitors.  In some cases, the distortion was quite low.  We have created
new designs for the sample holders for the JTO-funded experiments, and they were submitted for fabrication.  We also are
having samples prepared in advance of the tests.  New air nozzles were purchased.  Thermocouples are being purchased, along
with a data logger.

Operations/Commissioning:
The IR Demo was operated for two days this week to analyze phase noise in the drive laser and FEL output.
 

---Fred/dmag